\underline{1}

1

The Gnats and Gnus

A Document Preparation System^{∗}

Leslie A. Aamport^{†}

Abstract Starting from a historical point of view …

Keywords: Gnats, Gnus

1 Why Gnu

According to [?]^{1} …

• Some items are – bullets.

1. And some items (a) are numbers.

• But some items are

gnus A sort of large aninals.

1.1 Everything is changing

Change 1: Even fonts are changing.

∗to Gnu

∗to Barbara

†to Peter

1 Note

The figure is typeset in \figurefont. The caption body is typeset in

\captionbodyfont. The header of the caption is typeset in

                    \captionheaderfont.

Fig. 1: The body of the caption.